OpenStudy (anonymous):

Please help! I need to know if i did it right! What is the value of x to the nearest tenth? (drawing) i used tan

OpenStudy (anonymous):



OpenStudy (anonymous):

tan(angle)=op/adj tan75 = x/22 so 22tan75=x is this what you did? I think that's how you do it.

OpenStudy (anonymous):

ya thats what i did an i got 45.1 but the choices are 5.7 21.2 30.3 82.1 so i dont get what i did wrong

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ive spent about 30 minutes trying to figure out what i did wrong

OpenStudy (anonymous):

Okay is your calculator set to radians or degrees? lol I have had this happen to me before

OpenStudy (anonymous):

degrees

OpenStudy (anonymous):

Are you sure because I just used a calculator and the answer is 82.1.

OpenStudy (anonymous):

Just do tan75 and tell me what you get

OpenStudy (anonymous):

3.73

OpenStudy (anonymous):

Okay multiply that by 22 lol